# MVTamperBench: Evaluating Robustness of Vision-Language Models

### v4

@@ 4.1.1 Models Analysis Based on Performance
Figure 8: F1 (overall) scores for tampering types across model categories. Low & Moderate-performing models perform slightly better on Masking, while high-performing models show consistent robustness.
Figure 9: Variance in F1 (overall) scores across tampering types for each model category.
We analyze variance in model performances across tampering types in Figure 9. High-performing models exhibit negligible variance, indicating consistent robustness across all tampering effects. This stability suggests that these models are well-equipped to handle both spatial & temporal disruptions introduced by tampering.
Low-performing models, display minimal variance across most tampering types, with the exception of Masking, which shows a slightly higher variance. This highlights a specific weakness in processing visual obfuscations, likely due to their reliance on static features rather than robust temporal reasoning. The minimal variance across other tampering types suggests that these models fail uniformly, regardless of the type of manipulation.
Moderate-performing models demonstrate the highest variance overall, particularly for tampering types such as Dropping, Repetition, and Substitution. This behavior indicates an inconsistency in their ability to adapt to tampering effects. While these models achieve a balanced performance across easier tampering types, temporal disruptions like Dropping, Repetition, and Substitution pose significant challenges, leading to occasional spikes in variance. This suggests that moderate-performing models, while more capable than low-performing counterparts, still struggle to maintain robustness across a diverse set of tampering scenarios. We further analyze individual model performances in Appendix 7.3.1
4.1.2 Models Analysis Based on Model Size
We categorize models based on their parameter sizes into small (<7B), medium (7B–26B), and large (>26B) groups (Figure 11). While Figure 12 shows that larger models generally achieve higher F1 (overall) scores, Figure 5 confirms no significant correlation between model size and tampering detection performance (Pearson correlation = 0.05). A closer examination of individual model trends across size categories reveals several noteworthy patterns discussed in Appendix 7.3.2
@@ 4.1.3 Analysis Across Video Task Types
MVTamperBench comprises 19 video task categories, each evaluated across five tampering techniques. Figure 10 highlights the F1 (overall) scores for each task, averaged across all tampering types and models. While certain tasks exhibit high detection F1 score, others remain significantly more challenging. Appendix 7.3.3 provides additional insights into the performance trends of each tampering type across different tasks.
Figure 11: Distribution of Models by Size; Categories: Small (<7B), Medium (7B–26B), and Large (>26B)
@@
We further discuss the Key Findings & Future Directions in Appendix 7.5 & 7.6 respectively.
5 Conclusion
We introduced MVTamperBench, a comprehensive benchmark for evaluating the robustness of Multimodal Large Language Models (MLLMs) against five key video tampering techniques—Dropping, Masking, Repetition, Rotation, and Substitution. Through systematic experiments on 19 video tasks involving 45 models across 15+ families, we observed pronounced variability in resilience. Notably, even LMMs exceeding 70B parameters suffer severe performance drops, whereas select small models (<7absent7<7< 7B) demonstrate unexpectedly strong tampering detection, illustrating that size alone does not ensure robustness.
Beyond revealing these vulnerabilities, MVTamperBench offers actionable insights for refining model architectures and training pipelines, highlighting gaps that must be addressed before these systems can be reliably deployed in safety-critical settings. Our open-source framework supports reproducible evaluations and community-driven extensions, enabling researchers to integrate additional datasets or adapt tampering methods for domains like clickbait detection, content moderation, and surveillance feeds analysis.
Looking ahead, we plan to expand MVTamperBench with new tampering types (e.g., noise injection, frame shuffling) and domain-specific scenarios (e.g., healthcare, surveillance). By illuminating the nuanced impacts of video manipulations and guiding innovation in robust LMM architectures, MVTamperBench lays a strong foundation for next-generation multimodal models capable of withstanding adversarial manipulations.
@@ 6 Limitation
MVTamperBench provides a robust framework for evaluating LMM resilience against video tampering, but there are limitations that present opportunities for future work.
First, while the benchmark evaluates five tampering techniques, expanding to additional manipulations to better capture subtle, and emerging techniques. Second,the dataset currently relies on limited datasets. Expanding the benchmark to incorporate videos from diverse sources, such as user-generated content, surveillance footage, or policy enforcement, would broaden its applicability and relevance to task-specific and domain-specific challenges. Third, while the current evaluation focuses on binary detection, future benchmarks could assess a model’s ability to classify the specific tampering type, providing deeper insights into its robustness. Finally, scalability to closed-source and extremely large-scale models (>100absent100>100> 100B parameters) remains a challenge due to computational and cost constraints.
Addressing these limitations will enable MVTamperBench to further advance tampering detection and resilience across a broader range of applications.

### v5

@@ 4.2.1 Analysis based on Performance
Figure 8: F1 scores for tampering types across model categories. Low & Moderate-performing models perform slightly better on Masking, while high-performing models show consistent robustness.
Figure 9: Variance in F1 scores across tampering types for each model category.
We analyze the variance in model performance between tampering types in Figure 9. High-performing models exhibit negligible variance, indicating consistent robustness across all tampering effects. This stability suggests that these models are well-equipped to handle both spatial & temporal disruptions introduced by tampering.
Low-performing models display minimal variance across most tampering types, with the exception of Masking, which shows a slightly higher variance. This highlights a specific weakness in processing visual obfuscations, likely due to their reliance on static features rather than robust temporal reasoning. The minimal variance across other tampering types suggests that these models fail uniformly, regardless of the type of manipulation.
Moderate-performing models demonstrate the highest variance overall, particularly for tampering types such as Dropping, Repetition, and Substitution. This behavior indicates an inconsistency in their ability to adapt to tampering effects. While these models achieve a balanced performance across easier tampering types, temporal disruptions like Dropping, Repetition, and Substitution pose significant challenges, leading to occasional spikes in variance. This suggests that moderate-performing models, while more capable than low-performing counterparts, still struggle to maintain robustness across a diverse set of tampering scenarios. We provide a detailed analysis of individual model performance in Appendix A.4.1, and present qualitative trends based on model architecture and training paradigms across all models studied in Appendix A.4.2, offering insights into the factors influencing performance.
4.2.2 Analysis based on Model Size
We categorize models based on their parameter sizes into small (<7B), medium (7B–26B), and large (>26B) groups (Figure 11). While Figure 12 shows that larger models generally achieve higher F1 (overall) scores, Figure 5 confirms no significant correlation between model size and tampering detection performance (Pearson correlation = 0.05). A closer examination of individual model trends across size categories reveals several noteworthy patterns discussed in Appendix A.4.3.
@@ 4.2.3 Analysis across Video Task Types
MVTamperBench comprises 19 video task categories, each evaluated across five tampering techniques. Figure 10 highlights the F1 (overall) scores for each task, averaged across all tampering types and models. While certain tasks exhibit high detection F1 score, others remain significantly more challenging. Appendix A.4.4 provides additional insights into the performance trends of each tampering type across different tasks.
Figure 11: Distribution of models grouped by size; Categories: Small (<7B), Medium (7B–26B), and Large (>26B).
@@
We further discuss our Key Findings & Future Directions in Appendix A.6 & A.7 respectively.
5 Concluding Remarks
Novel Benchmark.
We introduced MVTamperBench, a comprehensive benchmark for evaluating the robustness of Multimodal Large Language Models (MLLMs) against five key video tampering techniques—Dropping, Masking, Repetition, Rotation, and Substitution. Through systematic experiments on 19 video tasks involving 45 models across 15+ families, we observed pronounced variability in resilience. Notably, even MLLMs exceeding 70B parameters suffer severe performance drops, whereas select small models (<7absent7<7< 7B) demonstrate unexpectedly strong tampering detection, illustrating that size alone does not ensure robustness.
Impact on Research Community.
Beyond revealing these vulnerabilities, we believe that MVTamperBench offers actionable insights for refining model architectures and training pipelines, highlighting gaps that must be addressed before these systems can be reliably deployed in safety-critical settings. Our open-source framework will support reproducible evaluations and community-driven extensions, enabling researchers to integrate additional datasets or adapt tampering methods for domains like clickbait detection, content moderation, and surveillance feeds analysis.
Future Work.
Looking ahead, we plan to expand MVTamperBench with new tampering types (e.g., noise injection, frame shuffling) and domain-specific scenarios (e.g., healthcare, surveillance). By illuminating the nuanced impacts of video manipulations and guiding innovation in robust MLLM architectures, MVTamperBench lays a strong foundation for next-generation multimodal models capable of withstanding adversarial manipulations.
@@ 6 Limitation
MVTamperBench provides a robust framework for evaluating MLLM resilience against video tampering, but there are limitations that present opportunities for future work.
First, while the benchmark evaluates five tampering techniques, expanding to additional manipulations/modalities to better capture subtle, and emerging techniques. Second,the dataset currently relies on limited datasets. Expanding the benchmark to incorporate videos from diverse sources, such as user-generated content, surveillance footage, or policy enforcement, would broaden its applicability and relevance to task-specific and domain-specific challenges. Third, while the current evaluation focuses on binary detection, future benchmarks could assess a model’s ability to classify the specific tampering type, providing deeper insights into its robustness. Finally, scalability to closed-source and extremely large-scale models (>100absent100>100> 100B parameters) remains a challenge due to computational and cost constraints.
Addressing these limitations will enable MVTamperBench to further advance tampering detection and resilience in a broader range of applications.